# Dynamic Subgraph Matching via Cost-Model-based Vertex Dominance Embeddings (Technical Report)

### v2

@@ 8.4. The DSM Efficiency Evaluation
Figure 14. DAS3 construction time on real/synthetic graphs, compared with baseline methods.
@@
The DSM Scalability Test w.r.t. Dynamic Graph Size $\bm{|V(G_{D})|}$: Figure 12(e) tests the scalability of our DSM approach with different dynamic graph sizes, $|V(G_{D})|$, from $10K$ to $1M$, where default values are assigned to other parameters. A larger dynamic graph incurs more matching candidate vertices (and, in turn, candidate subgraphs). From this figure, the time cost of DSM increases linearly with the increase of graph size $|V(G_{D})|$, nonetheless, remains low (i.e., 0.4 $\sim$ 104.62 $ms$), which confirms the scalability of our DSM approach for large graph sizes.
@@ 8.5. DAS3 Synopsis Initialization Cost
The DAS3 Synopsis Initialization Cost on Real/Synthetic Graphs. We compare the DAS3 synopsis initialization cost of our DSM approach (including time costs of vertex dominance embedding generation and DAS3 construction over vertex dominance embeddings) with online query time over real/synthetic graphs, where parameters are set to default values. In Figure 13, for graph size from 3K to 3.77M, the overall offline pre-computation time varies from 19 $ms$ $\sim$ 23 $sec$. Specifically, the time costs of embedding generation and DAS3 construction are 11 $ms$ $\sim$ 17 $sec$, 8 $ms$ $\sim$ 6 $sec$, respectively. On the other hand, since the online query time includes the time cost of DAS3 and embedding update, the maintenance cost of our DAS3 is low.
(a) real-world graphs
(b) synthetic graphs
@@
Figure 15. DAS3 memory cost on real/synthetic graphs, compared with baseline methods.
(a) $Syn\text{-}Uni$
(b) $Syn\text{-}Gau$
(c) $Syn\text{-}Zipf$
Figure 16. The DSM pre-computation and query costs w.r.t data graph size $|V(G_{D})|$.
@@
The DAS3 Index Time/Space Cost on Real/Synthetic Graphs. Since GF directly enumerates matching answers over the original data graph without any auxiliary data structure (Kankanamge et al., 2017; Sun et al., 2022a), Figures 14 and 15 compare the construction time and storage cost of our DAS3 with the other 4 baselines, where parameters are set to their default values. From Figure 14, we can find that our DAS3 needs shorter construction time over some complex graphs, i.e., ye, db, yt, and $Syn\text{-}Zipf$. This is because, different from baseline methods that build auxiliary data structure over initial query answers, our DAS3 is built over the initial data graph without enumerating the initial query answers and only depends on the data graph size. As shown in Figure 15, except SJ with exponential space cost (Choudhury et al., 2015; Sun et al., 2022a), our DAS3 needs more storage costs for storing/indexing vertex dominance embeddings. However, different from baseline methods that construct an index for each query graph during the online dynamic subgraph matching process, our DAS3 construction is offline and one-time only. Thus, our offline constructed DAS3 can be used to accelerate numerous online dynamic subgraph matching requests from users simultaneously with high throughput.
The DAS3 Synopsis Initialization Cost and Online Query Costs w.r.t. Data Graph Size $|V(G_{D})|$. Figure 16(a) evaluates the DAS3 synopsis initialization cost of our DSM approach, including time costs of the vertex dominance embedding generation and DAS3 construction over embeddings, compared with online DSM query time, over synthetic graphs $Syn\text{-}Uni$, where we vary the graph size $|V(G)|$ from $10K$ to $1M$ and other parameters are set to default values. Specifically, for graph sizes from $10K$ to $1M$, the time costs of the embedding generation and DAS3 construction are 0.03$\sim$3.11 $sec$ and 0.01$\sim$1.22 $sec$, respectively. The overall offline pre-computation time varies from 0.04 $sec$ to 4.33 $sec$, and the dynamic subgraph matching query cost is much smaller (i.e., 0.4$\sim$57.49 $ms$).
@@
In Figures 16(b) and 16(c), we can see similar experimental results over $Syn\text{-}Gau$ and $Syn\text{-}Zipf$ graphs, respectively, where the subgraph matching query cost is much smaller than the overall offline pre-computation time.
@@ 9. Related work
Dynamic Subgraph Matching: Existing works on exact dynamic subgraph matching (Fan et al., 2013b; Choudhury et al., 2015; Kankanamge et al., 2017; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021) can be classified into three categories: the recomputation-based (Fan et al., 2013b), direct-incremental (Kankanamge et al., 2017), and index-based incremental algorithms (Choudhury et al., 2015; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021). For the recomputation-based category, IncIsoMatch (Fan et al., 2013b) extracted a snapshot subgraph $g_{S}\subseteq G_{t}$ surrounding the updated edges at the timestamp $t$, conducted the subgraph matching over this snapshot subgraph $g_{S}$ (w.r.t. query graph $q$), and updated the answer set $A(q,t-1)$ at timestamp $(t-1)$ with new subgraph matching answers by removing redundant subgraphs. For the direct-incremental approach, GraphFlow (Kankanamge et al., 2017) computed incremental results over the dynamic graph by using a multi-way join from the updated edge.
Moreover, index-based incremental methods construct an index over the initial subgraph matching answers, and obtain incremental results from the index (instead of the data graph). For example, SJ-Tree (Choudhury et al., 2015) modeled a dynamic subgraph query as a multi-way join and organized all partial join results in an index to serve the query. TurboFlux (Kim et al., 2018) built a tree index in which each node contains candidates of a query vertex, and dynamically maintained the index to keep the consistency with each snapshot of the dynamic graph. SymBi (Min et al., 2021) improved the pruning power by constructing a graph index and designing an adaptive ordering method. IEDyn (Idris et al., 2017, 2020) maintained an index for candidate vertex sets and their connection edges, which can achieve constant delay enumeration for acyclic queries with graph homomorphism. In contrast, our work designs a novel and effective vertex dominance embedding technique for candidate vertex/subgraph retrieval, which can effectively prune vertex/subgraph candidates and improve query efficiency.
@@
In addition, there are also some studies that consider specific query graph topologies, such as paths (Peng et al., 2019; Sun et al., 2021), cycles (Qiu et al., 2018), and cliques (Mondal and Deshpande, 2016). Due to the high computation cost of finding exact matching results, several approximate algorithms (Chen and Wang, 2010; Fan et al., 2013b, 2010; Henzinger et al., 1995; Song et al., 2014) have been proposed. Different from these works, our paper focuses on the exact subgraph matching query over the dynamic graph, given general query graphs (i.e., with arbitrary query graph structures).
Graph Embeddings: Prior works on graph embedding usually transform graphs or nodes to $d$-dimensional embedding vectors for different downstream tasks, such as graph classification, vertex classification, link prediction, and so on. Several graph embedding methods used heuristic rules to encode a graph or its nodes, such as DeepWalk (Perozzi et al., 2014), LINE (Tang et al., 2015), SDNE (Wang et al., 2016), Node2vec (Grover and Leskovec, 2016), and Struc2vec (Ribeiro et al., 2017). These methods are designed for static graphs, which generate graph or node embeddings for fixed graph structures. Thus, we cannot directly use them for a dynamic graph task.
@@
Recently, the graph representation learning via graph neural networks (GNNs) has exhibited success in various graph-related applications (Sun et al., 2022b, 2024; Wang et al., 2020; Hao et al., 2021; Wang et al., 2021, 2022; Huang et al., 2022). With different network architectures, some previous works (Li et al., 2019; Bai et al., 2019; Duong et al., 2021; Ye et al., 2024) proposed to use GNNs to generate graph embeddings for the graph matching. However, these works either cannot guarantee the accuracy of tasks over unseen test graph data (Li et al., 2019; Bai et al., 2019) (due to the limitation of neural networks), or cannot efficiently and incrementally maintain embeddings in dynamic graphs with continuous updates (Duong et al., 2021; Ye et al., 2024). In contrast, our vertex dominance embedding technique does not use learning-based graph embedding, which can guarantee exact subgraph matching over dynamic graphs without false dismissals and enable incremental embedding maintenance over dynamic graphs upon updates.
10. Conclusions

### v3

@@ 8.4. The DIVINE Efficiency Evaluation
(a) The DIVINE efficiency on edge-insertion-only real/synthetic graphs, compared with baseline methods.
The DIVINE Efficiency on Edge-Insertion-Only Real/Synthetic Graphs: Figure 24(a) compares the efficiency of our DIVINE approach with that of 5 state-of-the-art baseline methods by varying the insertion ratio from 10% to 50% over both real and synthetic graphs, where all other parameters are set to default values. For GNN-PE over static graphs (Ye et al., 2024), due to the extremely high cost of GNN re-training, we cannot run the entire edge update stream, thus, we estimate the expected time cost for a period of consecutive (sampled) timestamps (with GNN re-training at least once). In Figures 8.4 and 8.4 with the default insertion ratio 10%, we can see that our DIVINE approach outperforms baseline methods mostly by 1-5 orders of magnitude. Note that DIVINE and all baselines require a higher time cost on Reddit (re) than other real/synthetic graphs, due to its high degree (i.e., avg_deg(G)=491.98avg\_deg(G)=491.98). Moreover, DIVINE performs better than GNN-PE by 3-5 orders of magnitude, since even one-time GNN re-training in GNN-PE incurs a much higher time cost than our DIVINE approach.
For other subfigures with different insertion ratios from 20% to 50%, we can see similar experimental results over both real and synthetic graphs. Although the total query time will increase as the number of inserted edges increases, our DIVINE approach can achieve performance that is up to 1-5 orders of magnitude higher than baseline methods.
real-world graphs ($Del_{10\%}$)] Refer to caption
real-world graphs ($Del_{20\%}$)] Refer to caption
real-world graphs ($Del_{30\%}$)] Refer to caption
real-world graphs ($Del_{40\%}$)] Refer to caption
real-world graphs ($Del_{50\%}$)] Refer to caption
synthetic graphs ($Del_{10\%}$)] Refer to caption
synthetic graphs ($Del_{20\%}$)] Refer to caption
synthetic graphs ($Del_{30\%}$)] Refer to caption
synthetic graphs ($Del_{40\%}$)] Refer to caption
synthetic graphs ($Del_{50\%}$)] Refer to caption
(b) The DIVINE efficiency on edge-deletion-only real/synthetic graphs, compared with baseline methods.
The DIVINE Efficiency on Edge-Deletion-Only Real/Synthetic Graphs: Since SJ and IED baselines do not support the edge deletion (Sun et al., 2022a), Figure 24(b) only compares the efficiency of our DIVINE approach with that of GF, TF, SYM, and GNN-PE over real/synthetic graphs, where the deletion ratio varies from 10% to 50% and all other parameters are set to their default values. From Figures 8.4 and 8.4 with the default deletion ratio 10%, our DIVINE approach can always outperform the four baseline methods. Overall, for all real/synthetic data sets (even for $re$ with 5.73MM edge deletions), the wall clock time of our DIVINE approach remains the lowest, compared with baselines.
From other subfigures with different deletion ratios 20% $\sim$ 50%, we can find similar experimental results over both real and synthetic graphs, where our DIVINE approach can always outperform the four baseline methods by up to 1-5 orders of magnitude.
real-world graphs ($Ins_{25\%}$)] Refer to caption
real-world graphs ($Ins_{50\%}$)] Refer to caption
real-world graphs ($Del_{25\%}$)] Refer to caption
real-world graphs ($Del_{50\%}$)] Refer to caption
real-world graphs ($Ins_{300K}$)] Refer to caption
synthetic graphs ($Ins_{25\%}$)] Refer to caption
synthetic graphs ($Ins_{50\%}$)] Refer to caption
synthetic graphs ($Del_{25\%}$)] Refer to caption
synthetic graphs ($Del_{50\%}$)] Refer to caption
synthetic graphs ($Ins_{300K}$)] Refer to caption
(c) The DIVINE efficiency with high/bursty edge update ratios, compared with baseline methods.
The DIVINE Efficiency with Different Edge Update Ratios: Figure 24(c) shows the efficiency of our DIVINE approach and baselines over real/synthetic graphs, for high edge insertion rates ($Ins$): 25% and 50%, high edge deletion rates ($Del$): 25% and 50%, and high/bursty updates of $300K$ edge insertions ($Ins_{300K}$), where other parameters are set to default values. Similar to results with default 10% edge insertion and deletion rates, respectively, our DIVINE approach always outperforms baselines by up to 1-5 orders of magnitude, for high or bursty edge update ratios.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(d) The DIVINE efficiency on real/synthetic graphs with vertex label changes, compared with baseline methods.
The DIVINE Efficiency on Dynamic Graphs with Vertex Label Changes: Figure 24(d) evaluates our DIVINE approach over real and synthetic graphs with vertex label changes, where we change 1% vertex labels upon every $10K$ updates, and default values are used for other parameters. As discussed in Section 6.3, our proposed DIVINE framework can naturally support dynamic vertex label changes by updating the embeddings of the vertex and its 1-hop neighbors. Similar to previous results (as shown in Figures 24(a) and 24(b)), our DIVINE approach outperforms the baseline methods for all real/synthetic graphs.
real-world graphs ($A_{1}$ mode)] Refer to caption
synthetic graphs ($A_{1}$ mode)] Refer to caption
real-world graphs ($A_{2}$ mode)] Refer to caption
synthetic graphs ($A_{2}$ mode)] Refer to caption
(e) The DIVINE efficiency with adversarial queries ($A_{1}$ and $A_{2}$ modes), compared with baseline methods.
The DIVINE Efficiency with Adversarial Queries ($\bm{A_{1}}$ and $\bm{A_{2}}$ Modes): Figure 24(e) shows our DIVINE query performance with adversarial queries over real/synthetic graphs, which have two modes, $A_{1}$” and $A_{2}$”. The $A_{1}$” mode tests 100 query graphs that do not occur in the data graph, whereas the $A_{2}$” mode evaluates 100 query graphs that lead to large candidate vertex sets. From subfigures, we can see that our DIVINE approach can always outperform baseline methods under both modes.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(f) The DIVINE efficiency with concurrent queries.
The DIVINE Efficiency with Concurrent Queries: Figure 24(f) illustrates the performance of our DIVINE approach, by using either a single thread (denoted as “s”) or multiple (16) threads in parallel (denoted as “m”), for 50 or 200 concurrent queries, which leads to four scenarios, “s50”, “m50”, “s200”, and “m200”, where default values are used for other parameters. Here, each thread is responsible for processing a query graph, including query embedding generation, synopsis search, refinement, and answer set update, such that CSM queries are independently answered in parallel. From subfigures, we can see that concurrent CSM queries can achieve lower time cost with multithreading for our DIVINE approach (instead of using a single thread), confirming the effectiveness of our DAS3 synopses to support parallel processing.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(g) The DIVINE efficiency of per-edge update, compared with baseline methods.
The DIVINE Efficiency of Per-Edge Update: Figure 24(g) illustrates our DIVINE per-edge update (insertion) cost, compared with baselines, where all parameters are set by default. From subfigures, we can observe that the per-edge update cost of our DIVINE approach is consistently lower than that of baselines over real/synthetic graphs, and remains low (i.e., 0.13 $\sim$ 89.11 $\mu s$).
To evaluate the robustness of our DIVINE approach, in the sequel, we test different scenarios, such as dynamic graph with edge deletions, vertex label changes, adversarial queries, and concurrent queries. We also vary parameter values on synthetic graphs (e.g., $|\Sigma|$, avg_deg(q)avg\_deg(q), $|V(q)|$, $avg\_deg(G_{D})$, and $|V(G_{D})|$), and omit the results of baselines to better illustrate the trends of curves.
DIVINE vs. $|\Sigma|$] Refer to caption
DIVINE vs. avg_deg(q)avg\_deg(q)] Refer to caption
DIVINE vs. $|V(q)|$] Refer to caption
DIVINE vs. $avg\_deg(G_{D})$] Refer to caption
DIVINE vs. $|V(G_{D})|$] Refer to caption
(h) The DIVINE efficiency w.r.t. parameters $|\Sigma|$, avg_deg(q)avg\_deg(q), $|V(q)|$, $avg\_deg(G_{D})$, and $|V(G_{D})|$.
[…]
The DIVINE Efficiency w.r.t. Average Degree, $\bm{avg\_deg(q)}$, of the Query Graph $\bm{q}$: Figure 8.4 examines the DIVINE performance by varying the average degree, avg_deg(q)avg\_deg(q), of the query graph $q$ from 2 to 4, where other parameters are set to default values. Higher degree avg_deg(q)avg\_deg(q) of $q$ incurs higher pruning power of query vertices. Therefore, when avg_deg(q)avg\_deg(q) increases, the DIVINE time cost decreases. For different avg_deg(q)avg\_deg(q) values, the DIVINE query cost remains low (i.e., 1.58 $\sim$ 7.68 $ms$).
The DIVINE Efficiency w.r.t. Query Graph Size $\bm{|V(q)|}$: Figure 8.4 illustrates our DIVINE performance, by varying the query graph size, $|V(q)|$, from 5 to 12, where default values are used for other parameters. When the number, $|V(q)|$, of vertices in query graph $q$ increases, fewer candidate subgraphs are expected to match with larger query graph $q$. On the other hand, larger query graph size $|V(q)|$ will cause higher query costs in finding candidates for more query vertices, through synopsis traversal and refinement. Therefore, the query time is influenced by these two factors. Nevertheless, the time cost remains low for different query graph sizes (i.e., <<5.365.36msms).
The DIVINE Efficiency w.r.t. Avg. Degree, $\bm{avg\_deg(G_{D})}$, of Dynamic Graph $\bm{G_{D}}$: Figure 8.4 presents our DIVINE performance with different average degrees, avg_degavg\_deg $(G_{D})$, of dynamic graph $G_{D}$, where $avg\_deg(G_{D})=3\sim 7$, and default values are used for other parameters. Intuitively, a higher degree $avg\_deg(G_{D})$ in the data graph $G_{D}$ incurs lower pruning power and more candidate vertices. Thus, when $avg\_deg(G_{D})$ becomes higher, the wall clock time also increases, especially for $Syn\text{-}Zipf$ (due to its skewed vertex label distribution). Nevertheless, the DIVINE query time remains small for different $avg\_deg(G_{D})$ values (i.e., 1.08 $\sim$ 5.82 $ms$).
@@
The DIVINE Scalability Test w.r.t. Dynamic Graph Size $\bm{|V(G_{D})|}$: Figure 24(h) tests the scalability of our DIVINE approach with different dynamic graph sizes, $|V(G_{D})|$, from $10K$ to $10M$, where default parameter values are used. A larger dynamic graph incurs more matching candidate vertices (and, in turn, candidate subgraphs). Thus, the time cost of our DIVINE approach increases with the increase of graph size $|V(G_{D})|$, nonetheless, remains the lowest (i.e., 0.4 $\sim$ 1.05 secsec for graph sizes up to $10M$), which confirms the scalability of our DIVINE approach for large graph sizes.
@@ 8.5. DAS3 Synopsis Initialization Cost
In this subsection, we report the memory cost of our DIVINE approach over real/synthetic graphs.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(i) The comparison analysis of the DIVINE offline pre-computation and online query costs on real/synthetic graphs.
@@
The DAS3 Synopsis Initialization Cost on Real/Synthetic Graphs. We compare the DAS3 synopsis initialization cost of our DIVINE approach (including time costs of vertex dominance embedding generation and DAS3 construction over vertex dominance embeddings) with online query time over real/synthetic graphs, where parameters are set to default values. In Figure 24(i), for graph sizes from 3K to 3.77M, the overall offline pre-computation time varies from 19 msms $\sim$ 147.52 secsec. Specifically, the time costs of embedding generation and DAS3 construction are 11 msms $\sim$ 141.85 secsec, 8 msms $\sim$ 5.67 secsec, respectively. On the other hand, since the online query time includes the time cost of DAS3 and embedding update, the maintenance cost of our DAS3 is low.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
@@
(j) DAS3 construction time on real/synthetic graphs, compared with baseline methods.
@@
The DAS3 Index Time Cost on Real/Synthetic Graphs. Since GF directly enumerates matching answers over the original data graph without any auxiliary data structure (Kankanamge et al., 2017; Sun et al., 2022a), Figure 24(j) compares the one-time construction cost of our DAS3 synopses with baselines, where parameters are set to their default values. From the figure, we can see that the time cost of our DAS3 synopses is comparable to that of baselines. Moreover, unlike baseline methods that construct an index over increasingly incoming (registered) query graphs $q$ (which may not be scalable), our DAS3 construction is one-time only over the initial data graph $G_{0}$. Thus, our DAS3 synopses can be used to accelerate numerous online CSM query requests from users simultaneously with high throughput.
real-world graphs] Refer to caption
synthetic graphs] Refer to caption
(k) The DAS3 memory cost on real/synthetic graphs, compared with baseline methods.
The DAS3 Space Cost on Real/Synthetic Graphs. Since GF directly enumerates matching answers over raw data graph without any auxiliary data structure (Kankanamge et al., 2017; Sun et al., 2022a), Figure 24(k) compares the storage cost of our DAS3 with that of the other 5 baselines, where parameters are set to default values. Our DIVINE approach always has lower space consumption of DAS3 synopses than GNN-PE over real/synthetic graphs, as well as SJ (Choudhury et al., 2015; Sun et al., 2022a) on most graphs. Moreover, DIVINE needs a bit higher space cost than other baselines. Nevertheless, different from these baselines that construct an index over more and more incoming (registered) query graphs $q$ (which may not be scalable), our DAS3 construction is one-time only over the initial graph $G_{0}$. Thus, our DAS3 synopses can be used to accelerate numerous online CSM query requests from users simultaneously with high throughput.
$Syn\text{-}Uni$] Refer to caption
$Syn\text{-}Gau$] Refer to caption
$Syn\text{-}Zipf$] Refer to caption
(l) The DIVINE pre-computation and query costs w.r.t data graph size $|V(G_{D})|$.
[…]
@@
In Figures 8.5 and 24(l), we can see similar experimental results over $Syn\text{-}Gau$ and $Syn\text{-}Zipf$ graphs, respectively, where the subgraph matching query cost is much smaller than the overall offline pre-computation time.
@@ 9. Related Work
Continuous Subgraph Matching: Due to the NP completeness of the subgraph isomorphism (Garey and Johnson, 1983; Cordella et al., 2004; Grohe and Schweitzer, 2020), continuous subgraph matching (CSM) is not tractable. Several exact algorithms over dynamic graphs have been proposed (Fan et al., 2013b; Choudhury et al., 2015; Kankanamge et al., 2017; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021), which can be classified into 3 categories: recomputation-based (Fan et al., 2013b), direct-incremental (Kankanamge et al., 2017), and index-based incremental algorithms (Choudhury et al., 2015; Idris et al., 2017, 2020; Kim et al., 2018; Min et al., 2021). The recomputation-based algorithms re-compute all the subgraph matching answers at each timestamp from scratch, the direct-incremental algorithms incrementally calculate new matching results from dynamic graphs upon updates, and index-based incremental ones incrementally maintain matching results via auxiliary indexes built over query answers.
Our proposed DIVINE approach falls into the second category of the direct-incremental algorithm. As shown in Algorithm 4, for already registered query graphs, our method incrementally maintains the answer set by starting from the updated edge and leveraging vertex dominance embeddings to efficiently identify affected candidates. This update procedure is executed on-the-fly and does not rely on any query-specific or materialized index structure, which aligns with the characteristics of a direct-incremental approach. However, unlike existing works that directly use structural information (e.g., vertex label and neighbors’ label set) to filter out vertex candidates, our work designs a novel and effective vertex dominance embedding technique for candidate vertex/subgraph retrieval, which can effectively prune vertex/subgraph candidates and improve query efficiency.
Furthermore, there exist graph-based SPARQL engines such as gStore (Zou et al., 2014), which transform RDF graphs and SPARQL queries to labeled data and query graphs, respectively, and apply a signature-based filtering technique followed by the subgraph isomorphism verification. Although gStore supports online updates over RDF repositories, it needs to recompute the matching answer after each update for dynamic query maintenance. In contrast, our DIVINE approach is designed for general-purpose, dynamic graphs, and leverages vertex dominance embeddings to support incremental subgraph matching with provable correctness and high efficiency.
@@
There are some studies on CSM w.r.t. specific query graph topologies (e.g., paths (Peng et al., 2019; Sun et al., 2021), cycles (Qiu et al., 2018), and cliques (Mondal and Deshpande, 2016)) or approximate matching (Chen and Wang, 2010; Fan et al., 2013b, 2010; Henzinger et al., 1995; Song et al., 2014). In contrast, we consider exact CSM over dynamic graphs, w.r.t. arbitrary query graph structures.
@@
Graph Embeddings: Traditional heuristic-based graph embedding methods (Perozzi et al., 2014; Tang et al., 2015; Wang et al., 2016; Grover and Leskovec, 2016; Ribeiro et al., 2017) are designed for static graphs and generate graph/node embeddings for fixed graph structures, which cannot be directly used for dynamic graphs. Some previous works (Li et al., 2019; Bai et al., 2019; Duong et al., 2021; Ye et al., 2024) proposed to use GNNs (Sun et al., 2022b, 2024; Wang et al., 2020; Hao et al., 2021; Wang et al., 2021, 2022; Huang et al., 2022) to generate graph embeddings for graph matching. However, these works either cannot guarantee the accuracy of tasks over unseen test graphs (Li et al., 2019; Bai et al., 2019) (due to limitations of neural networks), or cannot efficiently and incrementally maintain embeddings in dynamic graphs with continuous updates (Duong et al., 2021; Ye et al., 2024). In contrast, our proposed vertex dominance embeddings do not use learning-based graph embedding, which can guarantee exact CSM without false negatives and enable incremental embedding updates over dynamic graphs.
10. Conclusion